ribbon A is 1/3m long. it is 2/5m shorter than ribbon b. what is the total length of the ribbon?
Answer by dkppathak(439)

ribbon A is 1/3m long. it is 2/5m shorter than ribbon b. what is the total length of the ribbon?
given
A=1/3 m
as per given conditions
A=B-2/5 m or 1/3 =B-2/5
B=1/3+2/5=11/15m
total length of ribbon =A+B =1/3+11/15= 16/15 m
= 1 and 1/15 m